OpenStudy (sleepyjess):

Check 3 questions for me?

OpenStudy (sleepyjess):

Find 8u u = < -7, -2 > 8< -7, -2 > < -56, -16 >

OpenStudy (sleepyjess):

P = (5, 1) Q = (13, 5) Find the component form and magnitude of PQ Component form: 13 - 5, 8 - 1 (8, 7) Magnitude: \(\sf\sqrt{(v_1)^2+(v_2)^2}\\\) \(\sf\sqrt{8^2 + 7^2}\\\sqrt{64+49}\\\\\sqrt{113}\)

OpenStudy (sleepyjess):

U = < 4, -8 > V = < -2, 6 > Find u + v < 4 - 2, 8 + 6 > < 2, 14 >

OpenStudy (sleepyjess):

@YanaSidlinskiy

OpenStudy (sleepyjess):

@Directrix

OpenStudy (sleepyjess):

@perl will you check these?

OpenStudy (perl):

ok one moment

OpenStudy (perl):

first one is correct, 8u

OpenStudy (sleepyjess):

Ok, thanks :)

OpenStudy (perl):

P = (5, 1) Q = (13, 5) Find the component form and magnitude of PQ PQ = Q - P = (13,5) - (5,1) = ( 13 - 5 , 5 -1 ) so there is a small mistake < 8, 4> should be PQ

OpenStudy (sleepyjess):

Oh wow, thats for checking that, I have no idea where I got 8 from :P